Grade 11Algebra

x^4+1/X^4=47 find x+1/x what was answer in this question

krishna
7 Years agoGrade 11

1 Answer

Arun
7 Years ago
X⁴+1/x⁴=47
or, x⁴+1/x⁴+2-2=47
or, {(x²)²+2.x².1/x²+(1/x²)²}=47+2
or, (x²+1/x²)²=49
or, (x²+1/x²)²=7²
or, x²+1/x²=7 [neglecting the negative sign]
or, x²+1/x²+2-2=7
or, x²+2.x².1/x²+1/x²=7+2
or, (x+1/x)²=9
or, (x+1/x)²=3²
or, x+1/x=3 [negnecting the negative sign]
∴, x³+1/x³
=(x+1/x)³-3.x.1/x(x+1/x)
=(3)³-3(3)
=27-9
=18
 
hence (x + 1/x)^3 = 18 + 3 (x + 1/x)
 
t^3 – 3t – 18 = 0
 
t = 3
 
hence x + 1/x = 3